# Bridging Quantum Computing and Nuclear Structure: Atomic Nuclei on a Trapped-Ion Quantum Computer

### v2

@@ I Introduction
On near-term devices, variational approaches such as the Variational Quantum Eigensolver [2, 3] rely crucially on the choice of an ansatz that balances implementability and expressivity. Common families of ansatze — including unitary coupled cluster (UCC) variants — can be effective in some domains but often suffer from optimization difficulties, called barren plateaus [4], or unfavorable scaling of parameters and gate depth as system size grows. This situation has evoked a broader consideration of quantum algorithms beyond NISQ towards fault-tolerant quantum computing (FTQC), leading to a term like Early Fault-Tolerant Quantum Computing (E-FTQC) [5]. In such situations, it is essential to broadly examine each component of quantum computation—namely, state preparation (also simply called the ansatz), quantum algorithms, and measurement strategies for evaluating the observable of interest—for each specific problem.
@@ II.1 Target systems and Hamiltonian mapping
where $i$ denotes the single-particle state characterized by the quantum numbers $\left\{n,l,j,j_{z},t_{z}\right\}$. Here, $n$ is the principal quantum number, $l$ the orbital angular momentum, $j$ the total angular momentum, $j_{z}$ its projection of $j$ onto the $z$-axis, and $t_{z}$ the isospin projection.
@@
Thus, the pUCCD ansatz can be implemented as a sequence of Givens rotations on top of a reference state, e.g., initially prepared by applying $X$ gates to the qubits corresponding to occupied pair orbitals. Fig. 2 illustrates two circuit layouts for preparing the pUCCD ansatz, taking 22O nucleus as an example. (a) pUCCD(GS): optimized for devices with limited connectivity (e.g., IBM heavy-hexagon architectures), using Givens rotations followed by SWAP gates (GS) [21]. The basic idea behind the circuit is detailed in Ref. [38]. (b) pUCCD(G): one realization for all-to-all connectivity (e.g., trapped-ion devices) using only Givens rotations (G). In both cases, the number of parameters, angles of Givens rotations, is the same, given by the number of particle-hole pairs.
However, the circuit depth and number of two-qubit gates slightly differ. While pUCCD(GS) has a nearly unique circuit layout, pUCCD(G) admits many realizations depending on the choice of initial configuration and ordering of Givens rotations. We will therefore present simulator results for pUCCD(G) as distributions over multiple circuit realizations in Sec. III. In what follows, we use a single deterministic realization for both pUCCD(GS) and pUCCD(G) unless otherwise noted.
@@ III Results
Figure 3: Ground-state energies of target isotopes on noise-free statevector simulators. (a)–(c) Ground-state energies of oxygen, calcium, and nickel isotopes, respectively. Full configuration interaction (FCI, dotted lines with gray circles) and doubly occupied configuration interaction (DOCI, blue diamonds) are compared with statevector simulations using the pUCCD(GS) ansatz (green stars) and the pUCCD(G) ansatz (red triangles). Here, pUCCD(GS) denotes the pUCCD ansatz implemented with nearest-neighbor Givens rotations and SWAP gates, while pUCCD(G) assumes all-to-all connectivity, as realized in the Reimei device. (d)–(f) Differences between the pUCCD ansatz results and the DOCI results are shown relative to FCI in logarithmic scale. Uncertainty bands for pUCCD(G) are obtained by random sampling over circuits with different initial configurations and Givens-rotation orderings.
@@ III.1 Ideal simulation results of the pUCCD ansatz
Most importantly, pUCCD reproduces DOCI energies with an error of mostly less than $1\%$, while using far fewer parameters, demonstrating its compact yet accurate description. For reference, DOCI itself reproduces FCI ground-state energies within $0.1-1\%$, with deviations smaller than the typical uncertainties (a few percent) of modern nuclear many-body methods [47]. The residual differences depend on the target system: Ca isotopes in the $pf$ shell are well described within the zero-seniority space, whereas oxygen isotopes show larger contributions from higher-seniority configurations leading to slightly larger deviations from FCI. Nickel isotopes in the $jj45$ shell exhibit larger discrepancies between DOCI and pUCCD compared to oxygen and calcium isotopes. This arises from the factors intrinsic to the $jj45$ shell. The valence space, $jj45$ shell, consists of four $jj$-coupling orbitals ($1p_{1/2},1p_{3/2},0f_{5/2},0g_{9/2}$), and the configuration mixing among these orbitals is pronounced in Ni isotopes even for ground states since their effective single particle energies are close to each other [30]. The single-Trotter pUCCD ansatz, while efficient for capturing pairing correlations, becomes less expressive in the presence of such strong configuration mixing.
@@ III.2 Results on RIKEN-Quantinuum’s Reimei trapped-ion quantum device
As a whole, the hardware results agree closely with the noise-free simulations, showing relative errors mostly below $0.3\%$ to thedeal pUCCD results, and at most a few percent relative to FCI. These results demonstrate the capability of current quantum hardware to accurately capture nuclear pairing correlations.
@@
The accuracy on the diagonal term depends primarily on occupation numbers, while the $XX+YY$ terms are contribution is more sensitive to both occupation numbers and relative phases between configurations. As shown in the lower panel of Fig. 5, the distribution obtained using the Hadamard method is noticeably broader than that obtained with basis rotations.
@@ acknowledgments
This research was supported in part by JSPS Grant-in-Aid for Scientific Research (Grant Nos. JP22K14030, JP25H01511, JP25K01688), JST PRESTO (Grant No. JPMJPR25F8), JST COI-NEXT (Grant No. JPMJPF2221), MEXT Q-LEAP (Grant No. JPMXS0118067246), and RIKEN TRIP initiative (Nuclear Transmutation). This work is based on the results obtained from a project, JPNP20017, commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
@@ VI Uncertainty Bands of the pUCCD(G) ansatz
The fitted Gamma distributions were then used to construct the shaded uncertainty bands in Fig. (3) of the main text. We note that the pUCCD(GS) ansatz typically yields results close to the optimal edge of the pUCCD(G) distribution, except for cases such as 24O, 50Ca, and 52Ca where the qubit ordering or single-particle energy structure leads to less favorable initialization.

### v3

@@ I Introduction
On near-term devices, variational approaches such as the Variational Quantum Eigensolver [2, 3] rely crucially on the choice of an ansatz that balances implementability and expressivity. Common families of ansatze — including unitary coupled cluster (UCC) variants — can be effective in some domains but often suffer from optimization difficulties, called barren plateaus [4], or unfavorable scaling of parameters and gate depth as system size grows. This situation has evoked a broader consideration of quantum algorithms beyond noisy intermediate-scale quantum (NISQ) towards fault-tolerant quantum computing (FTQC), leading to a term like Early Fault-Tolerant Quantum Computing (E-FTQC) [5]. In such situations, it is essential to broadly examine each component of quantum computation—namely, state preparation (also simply called the ansatz), quantum algorithms, and measurement strategies for evaluating the observable of interest—for each specific problem.
@@ II.1 Target systems and Hamiltonian mapping
where $i$ denotes the single-particle state characterized by the quantum numbers $\left\{n,l,j,j_{z},t_{z}\right\}$. Here, $n$ is the principal quantum number, $l$ the orbital angular momentum, $j$ the total angular momentum, $j_{z}$ its projection onto the $z$-axis, and $t_{z}$ the isospin projection.
@@
Thus, the pUCCD ansatz can be implemented as a sequence of Givens rotations on top of a reference state, e.g., initially prepared by applying $X$ gates to the qubits corresponding to occupied pair orbitals. Fig. 2 illustrates two circuit layouts for preparing the pUCCD ansatz, taking 22O nucleus as an example. (a) pUCCD(GS): optimized for devices with limited connectivity (e.g., IBM heavy-hexagon architectures), using Givens rotations followed by SWAP gates (GS) [21]. The basic idea behind the circuit is detailed in Ref. [38]. (b) pUCCD(G): one realization for all-to-all connectivity (e.g., trapped-ion devices) using only Givens rotations (G). In both cases, the number of parameters is the same as the number of particle-hole pairs.
However, the circuit depth and number of two-qubit gates slightly differ. While pUCCD(GS) has a nearly unique circuit layout, pUCCD(G) admits many realizations depending on the choice of initial configuration and ordering of Givens rotations. We will therefore present simulator results for pUCCD(G) as distributions over multiple circuit realizations in Sec. III.
In what follows, we use a single deterministic realization for both pUCCD(GS) and pUCCD(G) unless otherwise noted. The qubits are aligned by descending order of $j$, total angular momentum of the corresponding single-particle states. This is not the optimal lowest-filling configuration; however, we verified that initializing qubits strictly in ascending SPE order is not necessarily superior. The present ordering is therefore a simple and reasonable choice among equivalent alternatives. One can further confirm that this choice does not give an outlier result by comparing with multiple random realizations, as will be shown later in Sec. III.
@@ III Results
Figure 3: Ground-state energies of target isotopes on noise-free statevector simulators. (a)–(c) Ground-state energies of oxygen, calcium, and nickel isotopes, respectively. Full configuration interaction (FCI, dotted lines with gray circles) and doubly occupied configuration interaction (DOCI, blue diamonds) are compared with statevector simulations using the pUCCD(GS) ansatz (green stars) and the pUCCD(G) ansatz (red triangles). (d)–(f) Differences between the pUCCD ansatz results and the DOCI results are shown relative to FCI in logarithmic scale. Uncertainty bands for pUCCD(G) are obtained by random sampling over circuits with different initial configurations and Givens-rotation orderings.
@@ III.1 Ideal simulation results of the pUCCD ansatz
Most importantly, pUCCD reproduces DOCI energies with an error of mostly less than $1\%$, while using far fewer parameters, demonstrating its compact yet accurate description. For reference, DOCI itself reproduces FCI ground-state energies within $0.1-1\%$, with deviations smaller than the typical uncertainties (a few percent) of modern nuclear many-body methods [47]. The residual differences depend on the target system: Ca isotopes in the $pf$ shell are well described within the zero-seniority space, whereas oxygen isotopes show larger contributions from higher-seniority configurations leading to slightly larger deviations from FCI. Nickel isotopes in the $jj45$ shell exhibit larger discrepancies between DOCI and pUCCD compared to oxygen and calcium isotopes. This arises from the factors intrinsic to the $jj45$ shell. The valence space of the $jj45$ shell consists of four $jj$-coupling orbitals ($1p_{1/2},1p_{3/2},0f_{5/2},0g_{9/2}$), and the configuration mixing among these orbitals is pronounced in Ni isotopes even for ground states since their effective single particle energies are close to each other [30]. The single-Trotter pUCCD ansatz, while efficient for capturing pairing correlations, becomes less expressive in the presence of such strong configuration mixing.
@@ III.2 Results on RIKEN-Quantinuum’s Reimei trapped-ion quantum device
As a whole, the hardware results agree closely with the noise-free simulations, showing relative errors mostly below $0.3\%$ to the ideal pUCCD results, and at most a few percent relative to FCI. These results demonstrate the capability of current quantum hardware to accurately capture nuclear pairing correlations.
@@
The accuracy on the diagonal term depends primarily on occupation numbers, while the $XX+YY$ terms are more sensitive to both occupation numbers and relative phases between configurations. As shown in the lower panel of Fig. 5, the distribution obtained using the Hadamard method is noticeably broader than that obtained with basis rotations.
@@ acknowledgments
This research was supported in part by JSPS Grant-in-Aid for Scientific Research (Grant Nos. JP22K14030, JP25H01511, JP25K01688), JST PRESTO (Grant No. JPMJPR25F8), JST ERATO (Grant No. JPMJER2304), JST COI-NEXT (Grant No. JPMJPF2221), MEXT Q-LEAP (Grant No. JPMXS0118067246), and RIKEN TRIP initiative (Nuclear Transmutation). This work is based on the results obtained from a project, JPNP20017, commissioned by the New Energy and Industrial Technology Development Organization (NEDO).
@@ VI Uncertainty Bands of the pUCCD(G) ansatz
The fitted Gamma distributions were then used to construct the shaded uncertainty bands in Fig. (3) of the main text. We note that the pUCCD(GS) ansatz typically yields results close to the optimal edge of the pUCCD(G) distribution, except for cases such as 24O, 50Ca, and 52Ca where the qubit ordering in ascending order of total angular momentum $j$ leads to less favorable initialization.